OpenStudy (anonymous):

log base 5 (5x+15y) expand the problem

OpenStudy (ranga):

\[\Large \log_{5} (5x+15y) = \log_{5} (5(x + 3y)) = \log_{5} 5 + \log_{5} (x+3y)\] \[\Large = 1 + \log_{5} (x+3y) \]

OpenStudy (anonymous):

thank you so much!

OpenStudy (ranga):

you are welcome.